# Non-Heisenbergian Quantum Mechanics

### v2

@@ Abstract
Relaxing the postulates of an axiomatic theory is a natural way to find more general theories, and historically, the discovery of non-Euclidean geometry is a famous example of this procedure. Here, we use this way to extend quantum mechanics by ignoring the heart of Heisenberg’s quantum mechanics—We do not assume the existence of a position operator that satisfies the Heisenberg commutation relation, $[\hat{x},\hat{p}]=i\hbar$. The remaining axioms of quantum theory, besides Galilean symmetry, lead to a more general quantum theory with a free parameter $l_{0}$ of length dimension, such that as $l_{0}\to 0$ the theory reduces to standard quantum theory. Perhaps surprisingly, this non-Heisenbergian quantum theory, without a priori assumption of the non-commutation relation, leads to a modified Heisenberg uncertainty relation, $\Delta x\Delta p\geq\sqrt{\hbar^{2}/4+l_{0}^{2}(\Delta p)^{2}}$, which ensures the existence of a minimal position uncertainty, $l_{0}$, as expected from various quantum gravity studies. By comparing the results of this framework with some observed data, which includes the first longitudinal normal modes of the bar gravitational wave detector AURIGA and the $2S-1S$ transition in the hydrogen atom, we obtain upper bounds on the $l_{0}$.
I Introduction
@@
There are various motivations to establish a theory of quantum mechanics without a position operator. Firstly, as shown by Pauli, there is no self-adjoint time operator conjugate to Hamiltonian operator Pauli (1958); Muga et al. (2007); Egusquiza and Muga (1999), and therefore, if one considers a self-adjoint position operator conjugate to momentum operator, space and time would not be on equal footing—a fact that is not in harmony with special relativity Toller (1999); Terno (2014); Caban et al. (2014), see also  [^1] In fact, there is no position operator that satisfies following natural demands Newton and Wigner (1949); Caban et al. (2014); Terno (2014): i) Its components be self-adjoint and commute with each other, ii) it satisfies canonical commutation relation with momentum operator, iii) It transforms like a three-vector under rotations, and iv) It transforms consistently under Lorentz boosts. The operator satisfying all of the conditions but the last condition is obtained by Newton and Wigner Newton and Wigner (1949): An eigen-state of Newton-Wigner position operator in one reference frame, which represents a localized state at a point, is not localized in other reference frames. Hence, the Newton-Wigner position operator have no simple covariant meaning Newton and Wigner (1949); Barr et al. (2023).. Secondly, and more importantly, as shown by Hegerfeldt, strictly localized states lead to faster-than-light probability propagation Hegerfeldt (1974); Hegerfeldt and Ruijsenaars (1980); Hegerfeldt (1985). This means that, in a self-consistent relativistic quantum mechanics, exactly localized states could not exist Thaller and Thaller (1984); Barat and Kimball (2003). Therefore, the position operator is not a well-defined concept if the eigenstates of the position operator represent exactly localized states. Although exactly localized states are not possible, unsharp localization is possible Bracken and Melloy (1999); Krekora et al. (2004, 2004); Kazemi et al. (2018), which can be described using a proper positive-operator-valued measure instead of a self-adjoint position operator Terno (2014); Kazemi et al. (2018). In fact, even in the non-relativistic setting, recent progress in the Wigner-Araki-Yanase theorem shows that exactly localized states are not practically accessible via position measurements Kuramochi and Tajima (2023); Busch and Loveridge (2011). Another motivation for avoiding exactly localized states comes from quantum gravity. Although a complete and agreed-upon theory of quantum gravity is still lacking, various approaches to quantum gravity Hossenfelder (2013) as well as gedankenexperiments that heuristically combine quantum theory and general relativity Calmet et al. (2004); Das and Vagenas (2008); Scardigli (1999); Adler and Santiago (1999); Amelino-Camelia suggest the existence of a limit to observable spatial resolution.
@@
The aim of the present work is to show how one can establish such a theory of quantum mechanics without a position operator. Although we avoid the position operator, (at least unsharp) position measurements are routinely performed in quantum experiments, even in the relativistic regime Sinha et al. (2010); Pleinert et al. (2021); Korzh et al. (2020). Therefore, a complete quantum theory must be able to predict the observed position distributions. In fact, a space-time discription of quamtum phenomena is critical for understanding the relativistic aspects of quantum information processes Caban et al. (2014); Peres and Terno (2004) and addressing some fundamental question in quantum theory Krekora et al. (2004); Wagner et al. (2011, 2012); Glasgow and Ware (2020); e.g., does photons satisfy uncertainty relations? If yes, how it could be proven without a reference to a position operator? Terno (2014); Bialynicki-Birula and Bialynicka-Birula (2012). Our starting point is the modern axiomatic formulation of (generalized) quantum theory in which the notion of observable is not limited to self-adjoint operators; an observable-distribution is given by a positive-operator-valued measure (POVM), not necessary a projection-valued measure (PVM) corresponding to a self-adjoint operator Masanes et al. (2019); Busch et al. (2009); Holevo (2011); Busch et al. (1989); Kunjwal et al. (2014). A justification for this generalization is provided by Gleason’s theorem Busch et al. (1996); Busch (2003), which guarantees that this structure is the most general one compatible with the probabilistic interpretation of quantum mechanics (for more details, see Busch et al. (1996)). Another justification, based on an operational approach to quantum theory, can be found in Masanes et al. (2019). In fact, the use of POVMs has been previously suggested to describe various observable distributions which their description via self-adjoint operators is problematic. Examples include the arrival time distribution Giannitrapani (1997); Hegerfeldt et al. (2010); Hegerfeldt and Muga (2010), work distribution Roncaglia et al. (2014), screen observable Werner (1986), and so on Busch et al. (2009). The aim of the present work is to establish a space-resolved quantum mechanics in which the position probability distribution is given by a proper position POVM at the fundamental level. Although the use of POVMs to describe particle localization has been previously suggested Busch et al. (1989), as far as we know, this idea has not been fully extended to systematically develop a complete quantum theory—including the derivation of its consequences on uncertainty relations, time evolution, the spectrum of interacting particle systems, and so on. Note that, in conventional quantum mechanics, the position operator not only determines the position probability density but also serves other roles, especially in canonical quantization where it is used to determine the Hamiltonian operator corresponding to a classical Hamiltonian. We will discuss how a proper position POVM could be consistently used for this purpose as well.
@@
In this work, our focus is on the non-relativistic regime; however, the results shed light on the subject in the general case. The relativistic generalization will be reported in future works. In this regard, the rest of the paper is organized as follows. In Section II, we obtain the general position probability density consistent with Galilean symmetry and the general structure of quantum theory. In Section III, we demonstrate that this general position probability density leads to a generalized position-momentum uncertainty relation, as well as a minimum position uncertainty, denoted as $l_{0}$. In Section IV, we discuss the time evolution in the presence of external potential and construct the associated Schrödinger equation without a position operator. Specifically, we study the simple harmonic oscillator in this framework and obtain an upper bound on $l_{0}$ by comparing the theoretical results with experimental data of the bar gravitational wave detector AURIGA. In Section V, we systematically generalize this formalism to interacting multi-particle systems, and compute the spectrum of the hydrogen atom in the presence of minimal position uncertainty. Using experimental data on $1S-2S$ transition, we derive a more stringent upper bound on $l_{0}$. In Section VI, we summarize the results and provide a discussion on future prospects.
II Probability density and Galilean symmetry
@@
According to generalized quantum mechanics Busch et al. (1989), we assume that, for any $\bm{x}\in\mathbb{R}^{3}$, there is an operator $\hat{Q}_{\bm{x}}$ such that the position probability density at $\bm{x}$ is given by
@@
In the conventional Heisenberg quantum mechanics, $\hat{Q}_{\bm{x}}$ is a projection operator constructed out of the eigenstates of a position operator, as $|\bm{x}\rangle\langle\bm{x}|$. We do not consider this assumption. Based on the probability theory, it is not necessary that $\hat{Q}_{\bm{x}}$ be a projection operator, it is only needed that $\hat{Q}_{\bm{x}}$ be a positive operator, $\hat{Q}_{\bm{x}}\geq 0$, and the set of $\hat{Q}_{\bm{x}}$ be normalized, i.e. $\int\hat{Q}_{\bm{x}}d^{3}x=\mathbb{1}$. These conditions make $\{\hat{Q}_{\bm{x}}|\bm{x}\!\in\!\mathbb{R}^{3}\}$ a POVM.
Now, let’s find the general position probability density for a free spinless particle consistent with non-relativistic space-time symmetries, including spatial and temporal translations, rotation, and Galilean boost. The space and time translation symmetries lead to $\hat{Q}^{t}_{\bm{x+a}}=e^{-i\bm{\hat{P}}.\bm{a}/\hbar}\hat{Q}^{t}_{\bm{x}}e^{i\bm{\hat{P}}.\bm{a}/\hbar}$, and $\hat{Q}^{t+\tau}_{\bm{x}}=e^{i\hat{H}\tau}\hat{Q}^{t}_{\bm{x}}e^{-i\hat{H}\tau/\hbar},$ respectively, which for infinitesimal transformations reduce to
@@
where $\hat{\bm{P}}/\hbar$ and $\hat{H}/\hbar$ are generators of space and time translation, respectively, and $\hat{Q}^{t}_{\bm{x}}$ is time-dependent position POVM in the Heisenberg picture, $\hat{Q}^{t}_{\bm{x}}\equiv e^{i\hat{H}t/\hbar}\hat{Q}_{\bm{x}}e^{-i\hat{H}t/\hbar}$. The equation (2) could be considered as an alternative for the usual position-momentum commutation relation. Using space and time translations symmetries, the $\hat{Q}^{t}_{\bm{x}}$ could be written in momentum representation as:
$\hat{Q}^{t}_{\bm{x}}\!=\!\!\int\!\!\ Q(\bm{p},\bm{k})e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}\left[e^{i\hat{H}t/\hbar}|\bm{k}\rangle\langle\bm{p}|e^{-i\hat{H}t/\hbar}\right]d^{3}\bm{p}d^{3}\bm{k},$
where, $Q(\bm{p},\bm{k})=\langle\bm{k}|\hat{Q}^{0}_{\bm{0}}|\bm{p}\rangle$. Similarly, the Galilean boost and rotation symmetries lead to the following commutation relations, respectively:
@@
Here, $\hat{\bm{K}}/\hbar$ and $\hat{\bm{J}}/\hbar$ are generators of the boost and rotation transformations in Hilbert space. Then, using the Lie algebra of the projective representation of the Galilean group Weinberg (1995); Ballentine (2014), the equations (5) and (4) lead to the following constraints on the $Q(\bm{p},\bm{k})$ function:
@@
which together ensure the $Q(\bm{p},\bm{k})$ is a function of $|\bm{p}-\bm{k}|$. Therefore, the general bilinear position probability density as (1) and consistent with non-relativistic space-time symmetries, Eq. (2-5), is given by
$\rho_{t}(\bm{x})\!=\!\!\!\!\int\!f(|\bm{p}-\bm{k}|)\tilde{\psi}_{t}(\bm{p})\tilde{\psi}_{t}^{*}(\bm{k})e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}d^{3}\bm{p}d^{3}\bm{k},$ (6)
where $\tilde{\psi}_{t}(\bm{p})$ is momentum space wave function, $\langle\bm{p}|\psi_{t}\rangle$, and $f(|\bm{p}-\bm{k}|)$ is an isotropic positive-definite kernel. The kernel $f(|\bm{p}-\bm{k}|)$ is also called a radial positive definite function. The characterization of radial positive definite functions is a classical result of I. Schoenberg Schoenberg (1938): A function $f(r)$ is radial positive definite on $\mathbb{R}^{n}$, If and only If there exists a finite positive measure $\alpha$ on $\mathbb{R}^{+}$ such that
[…]
where $\Omega_{n}(r)=\Gamma(n/2)(2/r)^{(n-2)/2}J_{(n-2)/2}(r),$ and $J_{n}$ is the Bessel function of order $n$ (see Wells and Williams (2012); Gneiting (1999); Wendland (2004) for further details). It is easy to see that the probability density which is given in (6) is conserved; it satisfies continuity equation, $\partial\rho/\partial t+\nabla.\bm{J}=0,$ where the probability current $\bm{J}$ is given by
$\bm{J}(\bm{x})=\int\frac{\bm{p}+\bm{k}}{2m}f(|\bm{p}-\bm{k}|)\tilde{\psi}(\bm{p})\tilde{\psi}^{*}(\bm{k})e^{i(\bm{p}-\bm{k}).\bm{x}/\hbar}d^{3}\bm{p}d^{3}\bm{k},$ (7)
for a free particle with $\hat{H}=\hat{\bm{P}}^{2}/2m$. Note that the only difference between the equations (6) and (7) and their counterparts in conventional quantum theory is the existence of the kernel $f(|\bm{p}-\bm{k}|)$.
III Modified uncertainty relation

### v4

@@ Abstract
Relaxing the postulates of an axiomatic theory is a natural way to find more general theories, and historically, the discovery of non-Euclidean geometry is a famous example of this procedure. Here, we use this way to extend quantum mechanics by ignoring the heart of Heisenberg’s quantum mechanics—We do not assume the existence of a position operator that satisfies the Heisenberg commutation relation, $\hat{x},\hat{p}]=i\hbar$. The remaining axioms of quantum theory, besides Galilean symmetry, lead to a more general quantum theory with a free parameter $l_{0}$ of length dimension, such that as $l_{0}\to 0$ the theory reduces to standard quantum theory. Perhaps surprisingly, this non-Heisenbergian quantum theory, without a priori assumption of the non-commutation relation, leads to a modified Heisenberg uncertainty relation, $\Delta x\Delta p\geq\sqrt{\hbar^{2}/4+l_{0}^{2}(\Delta p)^{2}}$, which ensures the existence of a minimal position uncertainty, $l_{0}$, as expected from various quantum gravity studies. By comparing the results of this framework with some observed data, which includes the first longitudinal normal modes of the bar gravitational wave detector AURIGA and the 2S1S2S-1S transition in the hydrogen atom, we obtain upper bounds on the $l_{0}$.
1 Introduction
@@
There are various motivations to establish a theory of quantum mechanics without a position operator. Firstly, as shown by Pauli, there is no self-adjoint time operator conjugate to the Hamiltonian operator [1, 2, 3], and therefore, if one considers a self-adjoint position operator conjugate to the momentum operator, space and time would not be on an equal footing—a fact that is not in harmony with special relativity [^1] In fact, there is no position operator that satisfies following natural demands [4, 5, 6]: i) Its components be self-adjoint and commute with each other, ii) it satisfies canonical commutation relation with momentum operator, iii) It transforms like a three-vector under rotations, and iv) It transforms consistently under Lorentz boosts. The operator satisfying all of the conditions but the last condition is obtained by Newton and Wigner [4]: An eigen-state of Newton-Wigner position operator in one reference frame, which represents a localized state at a point, is not localized in other reference frames. Hence, the Newton-Wigner position operator have no simple covariant meaning [4, 7].[8, 6, 5]. Secondly, and more importantly, as shown by Hegerfeldt, strictly localized states lead to faster-than-light probability propagation [9, 10, 11]. This means that, in a self-consistent relativistic quantum mechanics, exactly localized states cannot exist [12, 13]. Therefore, the position operator is not a well-defined concept if the eigenstates of the position operator represent exactly localized states. Although exactly localized states are not possible, unsharp localization is possible [14, 15, 16], which can be described using a proper positive-operator-valued measure instead of a self-adjoint position operator [6, 16]. In fact, even in the non-relativistic setting, recent progress regarding the Wigner-Araki-Yanase theorem shows that exactly localized states are not practically accessible via position measurements [17, 18]. Another motivation for avoiding exactly localized states comes from quantum gravity. Although a complete and agreed-upon theory of quantum gravity is still lacking, various approaches to quantum gravity [19], as well as gedankenexperiments that heuristically combine quantum theory and general relativity [20, 21, 22, 23, 24] suggest the existence of a limit to observable spatial resolution.
@@
The aim of the present work is to show how one can establish such a theory of quantum mechanics without a position operator. Although we avoid the position operator, (at least unsharp) position measurements are routinely performed in quantum experiments, even in the relativistic regime [25, 26, 27]. Therefore, a complete quantum theory must be able to predict the observed position distributions. In fact, a space-time description of quantum phenomena is critical for understanding the relativistic aspects of quantum information processes [5, 28] and addressing some fundamental questions in quantum theory [15, 29, 30, 31]. For example, do photons satisfy the position-momentum uncertainty relation? If so, how can it be proven without reference to a position operator [6, 32]? Our starting point is the modern axiomatic framework of (generalized) quantum theory, where the observables are not restricted to self-adjoint operators. Instead, in the general case, an observable is described by a positive-operator-valued measure (POVM), which is not necessarily a projection-valued measure (PVM) corresponding to a self-adjoint operator [33, 34, 35, 36, 37]. A justification for this generalization is provided by Gleason’s theorem [38, 39], which guarantees that this structure is the most general one compatible with the probabilistic interpretation of quantum mechanics (for more details, see [38]). Another justification, based on an operational approach to quantum theory, is given in [33]. In fact, the use of POVMs has been previously suggested to describe various observable distributions whose description via self-adjoint operators is problematic. Examples include the arrival time distribution [40, 41, 42], work distribution [43], screen observable [44], among others [34]. The aim of the present work is to establish a space-resolved quantum mechanics in which the position probability distribution is given by a proper position POVM at the fundamental level. Although the use of POVMs to describe particle localization has been previously suggested [36], to the best of our knowledge, this idea has not yet been fully extended to systematically develop a complete quantum theory—including deriving its consequences for uncertainty relations, time evolution, the spectrum of interacting particle systems, and so on. Note that in conventional quantum mechanics, the position operator not only determines the position probability density but also serves other roles, especially in canonical quantization where it is used to determine the Hamiltonian operator corresponding to a classical Hamiltonian. We will discuss how a proper position POVM can also be consistently used for this purpose.
@@
In this work, our focus is on the non-relativistic regime; however, the results shed light on the subject in the general case. The relativistic generalization will be reported in future works. In this regard, the rest of the paper is organized as follows. In Section II, we obtain the general position probability density consistent with Galilean symmetry and the general structure of quantum theory. In Section III, we show that this general position probability density gives rise to a generalized position-momentum uncertainty relation and implies a minimum position uncertainty, denoted by $l_{0}$. In Section IV, we discuss the time evolution in the presence of external potential and construct the associated Schrödinger equation without a position operator. Specifically, we study the simple harmonic oscillator in this framework and obtain an upper bound on $l_{0}$ by comparing the theoretical results with experimental data from the bar gravitational wave detector AURIGA. In Section V, we systematically generalize this formalism to interacting multi-particle systems, and compute the spectrum of the hydrogen atom in the presence of minimal position uncertainty. Using experimental data on the 1S2S1S-2S transition, we derive a more stringent upper bound on $l_{0}$. In Section VI, we summarize the results and provide a discussion on future prospects.
2 Probability density and Galilean symmetry
@@
According to generalized quantum mechanics [36], we assume that, for any $\bm{x}\in\mathbb{R}^{3}$, there is an operator $\hat{Q}_{\bm{x}}$ such that the position probability density at $\bm{x}$ is given by
@@
In the conventional Heisenberg quantum mechanics, $\hat{Q}_{\bm{x}}$ is a projection operator constructed from the eigenstates of a position operator, as $|\bm{x}\rangle\langle\bm{x}|$. We do not adopt this assumption. From the perspective of probability theory, it is not necessary for $\hat{Q}_{\bm{x}}$ to be a projection operator. Instead, it is only required that $\hat{Q}_{\bm{x}}$ be a positive operator, $\hat{Q}_{\bm{x}}\geq 0$, and that the set of $\hat{Q}_{\bm{x}}$ be normalized, i.e., $\int\hat{Q}_{\bm{x}}d^{3}x=\mathbb{1}$. These conditions make $\{\hat{Q}_{\bm{x}}|\bm{x}\!\in\!\mathbb{R}^{3}\}$ a POVM.
Now, let’s find the general position probability density for a free spinless particle consistent with non-relativistic space-time symmetries, including spatial and temporal translations, rotation, and Galilean boost. The space and time translation symmetries lead to $\hat{Q}^{t}_{\bm{x+a}}=e^{-i\bm{\hat{P}}.\bm{a}/\hbar}\hat{Q}^{t}_{\bm{x}}e^{i\bm{\hat{P}}.\bm{a}/\hbar}$ and $\hat{Q}^{t+\tau}_{\bm{x}}=e^{i\hat{H}\tau/\hbar}\hat{Q}^{t}_{\bm{x}}e^{-i\hat{H}\tau/\hbar},$ respectively. In the infinitesimal limit, these transformations yield
@@
where $\hat{\bm{P}}$$\hbar$ and $\hat{H}$$\hbar$ are generators of space and time translations, respectively, and $\hat{Q}^{t}_{\bm{x}}$ is the time-dependent position POVM in the Heisenberg picture, defined as $\hat{Q}^{t}_{\bm{x}}\equiv e^{i\hat{H}t/\hbar}\hat{Q}_{\bm{x}}e^{-i\hat{H}t/\hbar}$. Equation (2) encapsulates the infinitesimal translational symmetry in our framework, analogous to the canonical commutation relation in standard quantum theory [^2] Both Eq. (2) in our model and the position–momentum commutation relation in standard quantum mechanics originate from (infinitesimal) spatial translation symmetry [45, 46]. In fact, Eq. (2) leads to the standard commutation relation, $\hat{X}_{i},\hat{P}_{j}]=i\hbar\delta_{ij}$, provided that $\hat{\bm{X}}$ is defined as $\int\bm{x}\,\hat{Q}_{\bm{x}}\,d^{3}x$. However, $\hat{\bm{X}}$ coincides with the conventional position operator only when exactly localized states $|\bm{x}\rangle$ exist and $\hat{Q}_{\bm{x}}$ is defined as the projector $\hat{Q}_{\bm{x}}=|\bm{x}\rangle\langle\bm{x}|$. In the general case, when $\hat{Q}_{\bm{x}}$ is not a projector, $\hat{\bm{X}}$ cannot be interpreted as a position operator in the usual sense, since $\langle\bm{x}^{2}\rangle\equiv\int\rho(\bm{x})\,\bm{x}^{2}\,d^{3}x\neq\langle\psi|\hat{\bm{X}}^{2}|\psi\rangle$.. Using space and time translation symmetries, $\hat{Q}^{t}_{\bm{x}}$ could be written in momentum representation as:
$\hat{Q}^{t}_{\bm{x}}\!=\!\!\int\!\!\!\ Q(\bm{p},\bm{k})e^{\frac{i}{\hbar}(\bm{p}-\bm{k}).\bm{x}}\!\left[e^{i\frac{\hat{H}t}{\hbar}}|\bm{k}\rangle\langle\bm{p}|e^{-i\frac{\hat{H}t}{\hbar}}\right]\!d^{3}\bm{p}d^{3}\bm{k},$
where $Q(\bm{p},\bm{k})=\langle\bm{k}|\hat{Q}^{0}_{\bm{0}}|\bm{p}\rangle$. Similarly, the Galilean boost and rotation symmetries lead, respectively, to the following commutation relations:
@@
Here, $\hat{\bm{K}}$$\hbar$ and $\hat{\bm{J}}$$\hbar$ denote the generators of the boost and rotation transformations in Hilbert space. Then, using the Lie algebra of the projective representation of the Galilean group [47, 45], the equations (5) and (4) lead to the following constraints on the function $Q(\bm{p},\bm{k})$:
@@
which together ensure that $Q(\bm{p},\bm{k})$ is a function of $|\bm{p}-\bm{k}|$—The mathematical details of this derivation are provided in Appendix A. Therefore, the general bilinear position probability density consistent with the non-relativistic space-time symmetries expressed in Eqs. (2)–(5) is given by:
$\rho_{t}(\bm{x})\!=\!\!\int\!f(|\bm{p}-\bm{k}|)\tilde{\psi}_{t}(\bm{p})\tilde{\psi}_{t}^{*}(\bm{k})e^{\frac{i}{\hbar}(\bm{p}-\bm{k}).\bm{x}}d^{3}\bm{p}d^{3}\bm{k},$ (6)
where $\tilde{\psi}_{t}(\bm{p})$ is the momentum-space wave function, $\langle\bm{p}|\psi_{t}\rangle$, and $f(|\bm{p}-\bm{k}|)$ is an isotropic positive-definite kernel. The kernel $f(|\bm{p}-\bm{k}|)$ is also known as a radial positive definite function. The characterization of radial positive definite functions is a classical result of I. Schoenberg [48]: A function $f(r)$ is radial positive definite on $\mathbb{R}^{n}$ if and only if there exists a finite positive measure $\alpha$ on $\mathbb{R}^{+}$ such that
$f(r)=\int_{0}^{\infty}\Omega_{n}(ru)d\alpha(u),$
where $\Omega_{n}(r)=\Gamma(n/2)(2/r)^{(n-2)/2}J_{(n-2)/2}(r),$ and $J_{n}$ is the Bessel function of order $n$ (see [49, 50, 51] for further details).
It is straightforward to verify that the probability density given in (6) is conserved; it satisfies the continuity equation, $\partial\rho/\partial t+\nabla.\bm{J}=0,$ where the probability current $\bm{J}$ is given by
$\bm{J}\!=\!\!\int\!\frac{\bm{p}+\bm{k}}{2m}f(|\bm{p}-\bm{k}|)\tilde{\psi}(\bm{p})\tilde{\psi}^{*}(\bm{k})e^{\frac{i}{\hbar}(\bm{p}-\bm{k}).\bm{x}}d^{3}\bm{p}d^{3}\bm{k},$ (7)
for a free particle with $\hat{H}=\hat{\bm{P}}^{2}/2m$. It should be noted that the only difference between the equations (6) and (7) and their counterparts in conventional quantum theory is the appearance of the kernel $f(|\bm{p}-\bm{k}|)$.
3 Modified uncertainty relation